### v1

# Do We Need Large VLMs for Spotting Soccer Actions?

Abstract
Traditional video-based tasks like soccer action spotting rely heavily on visual inputs, often requiring complex and computationally expensive models to process dense video data. In this work, we propose a shift from this video-centric approach to a text-based task, making it lightweight and scalable by utilizing Large Language Models (LLMs) instead of Vision-Language Models (VLMs). We posit that expert commentary, which provides rich, fine-grained descriptions and contextual cues such as excitement and tactical insights, contains enough information to reliably spot key actions in a match. To demonstrate this, we use the SoccerNet Echoes dataset, which provides timestamped commentary, and employ a system of three LLMs acting as judges specializing in outcome, excitement, and tactics. Each LLM evaluates sliding windows of commentary to identify actions like goals, cards, and substitutions, generating accurate timestamps for these events. Our experiments show that this language-centric approach performs effectively in detecting critical match events, providing a lightweight and training-free alternative to traditional video-based methods for action spotting.
Do We Need Large VLMs for Spotting Soccer Actions?
Ritabrata Chakrabortythanks: Corresponding author Rajatsubhra Chakraborty Manipal University Jaipur UNC–Charlotte Avijit Dasgupta Sandeep Chaurasia IIIT Hyderabad Manipal University Jaipur ritabrata.229301716@muj.manipal.edu
1 Introduction
Football is a game of mistakes. Whoever makes the fewest mistakes wins.
Johan Cruyff
In the domain of video understanding, visual frames have traditionally been considered the best input for many tasks, including action spotting, event detection, and object recognition. However, these methods often require significant computational resources to process and analyze the dense video data. Despite the advancements in video models, such as convolutional neural networks (CNNs) and vision transformers (ViTs), the need for high-resolution video inputs can be prohibitive in both training and deployment scenarios. In many video contexts, particularly in sports, there is a spoken commentary that accompanies the visual content. This commentary provides detailed descriptions of the ongoing actions, events, and emotions, offering an alternative source of information. It is intriguing to consider whether these spoken descriptions, which can often provide rich contextual cues, could replace or complement visual inputs for specific tasks. By focusing on the textual information provided by commentators, we may gain a lightweight, scalable solution that avoids the complexity of video-based models.
Action spotting Seweryn et al. (2023), a core task in sports analytics, aims to identify key events within a video, such as goals, penalties, or substitutions, by analyzing the visual content. Traditional approaches have relied on object detection and tracking techniques that require parsing every frame of the video to detect specific actions. These methods can be computationally expensive and often struggle with long sequences or multiple simultaneous events. In contrast, when considering the commentary, each moment in the match is often described in rich detail, including the action, the players involved, and the contextual relevance. The spoken word can provide a nuanced understanding of the match dynamics, capturing moments of excitement, controversy, and strategic importance that may not always be fully conveyed through visual data alone. This raises an interesting possibility: Can we leverage textual commentary as a primary input for action spotting, bypassing the need for video frames?
In this paper, we explore this question by converting the traditional video-based action spotting task into a text-based one. Specifically, we investigate whether expert commentary, which provides fine-grained descriptions and contextual cues, is sufficient for reliably spotting key actions. The use of Large Language Models (LLMs) enables us to process and evaluate these textual descriptions, which contain detailed insights into match events, without requiring visual data. This approach not only alleviates the need for computationally intensive video models but also opens up new avenues for action detection in contexts where visual data may be scarce or of low quality, such as archived footage or radio broadcasts.
Figure 1: Our proposed LLM-based action spotting framework.
We define two key research questions (RQs) for our study:
Research Question 1 Can fine-grained commentary provide sufficient information for accurate action spotting? In this study, we investigate whether the detailed descriptions provided by commentators are enough to reliably detect key actions, such as goals, cards, and substitutions, without relying on visual data.
Research Question 2 How can we leverage Large Language Models (LLMs) for event detection in textual commentary? We examine how a system of three LLMs, each specialized in outcome, excitement, and tactics, can be used to evaluate timestamped commentary windows and identify event-worthy segments.
Through our study, we provide the following contributions:
  • We utilize the SoccerNet Echoes Gautam et al. (2024) dataset, which contains timestamped commentary of soccer matches, to reformulate action spotting as a text-based task.
  • We design and implement a three-LLM system that judges the commentary based on outcome (goals, cards), excitement (crowd reactions, dramatic moments), and tactics (substitutions, formations).
  • We demonstrate that fine-grained commentary, in many cases, provides comparable or even superior information for event detection compared to visual cues.
  • We show that, by focusing on commentary alone, it is possible to detect key events reliably, highlighting the potential of language-centric models for sports analytics.
The rest of the paper is structured as follows. Section 2 discusses present literature around using text-based inputs for video tasks and action spotting in soccer matches. Section 3 explains our proposed framework in detail. Section 4 sheds light on the experimental setup and quantitative results. Finally we discuss some limitations in Section 5.
2 Related Work
Fine-Grained Descriptions in Video-Based Tasks.
In video-based understanding tasks, traditional models have primarily relied on visual features extracted from video frames to detect and classify events. However, recent research has begun exploring the use of fine-grained descriptions—specifically, textual information derived from transcriptions or commentary—to enhance performance in tasks like action spotting and event detection. Xie et al. (2019) demonstrated that integrating visual information with text can improve performance in action recognition tasks, as descriptive cues often convey context that is missed by raw visual data. Fine-grained descriptions, which often come from sources such as narrators or commentators, offer a complementary modality that is particularly useful in contexts where visual cues are ambiguous or incomplete. In addition, Su et al. (2012) highlighted the utility of crowd-sourced commentary to aid in object detection tasks, which suggests that action spotting in dynamic environments, such as sports, could be enhanced by considering detailed narrative descriptions. These works point toward the potential of substituting or supplementing visual cues with textual descriptions, thereby providing more precise context for video interpretation tasks.
Action Spotting in Soccer Videos.
Action spotting in soccer has long relied on visual inputs, particularly tracking players and ball movements. However, recent developments in leveraging commentary and other textual sources for action detection have gained attention. Giannakopoulos et al. (2016) proposed a method that uses timestamped commentary as input to detect key moments in soccer, such as goals or penalties, demonstrating that textual data can complement traditional visual cues. Another approach by Andrews et al. (2024) used a multi-modal network that combines both video frames and textual commentary to detect key events in football matches. These methods, however, often require substantial computational resources and may struggle with temporal consistency between visual events and their descriptions. While these advancements are promising, they highlight the gap in efficiently using fine-grained commentary without relying heavily on computationally expensive visual inputs or needing large amounts of data for training.
Despite these promising advancements, there remains a gap in fully utilizing fine-grained commentary for video understanding tasks like action spotting, especially in the context of soccer. Existing methods either rely on computationally expensive visual cues or fail to achieve consistent performance with textual input alone. In this paper, we address this gap by proposing a novel method that employs a three-LLM (Large Language Model) judge setup for action spotting based purely on textual commentary.
3 Methodology
SoccerNet-Echoes Dataset.
The SoccerNet-Echoes dataset is an essential component of our approach. This dataset provides detailed English commentary for soccer matches, which is transcribed from video using Whisper-based Automatic Speech Recognition (ASR). Whisper Radford et al. (2022) is known for its ability to transcribe multiple languages and its robustness to different accents, making it suitable for real-time, dense captioning in soccer broadcasts. The transcriptions are then formatted as timestamped text, which is far more concise and contextually richer than the video-based dense captions often provided by traditional action spotting models. The commentary text provides rich contextual information that can guide the identification of meaningful game actions, such as goals, fouls, and tactical shifts, and is the primary input to our semantic judges.
Large Language Model Judges.
We use Llama 3.1 8B Grattafiori et al. (2024) to create three specialized semantic judges, each focusing on different aspects of the game, highlighted in Figure 1.Each judge is steered by a dedicated system prompt plus 2–3 few-shot exemplars.
[…]
Goals, penalties, red/yellow cards.
[…]
Crowd eruptions, near-misses, controversies.
[…]
Substitutions, formation changes, momentum swings.
Our classifier goes from binary to class-level matching, providing a quick look (Yes/No) for majority voting (next step), and a detailed classwise matching as highlighted below
Input: full English commentary for a 10 s window. Output: 1. YES/NO – does the window contain an action-worthy event? 2. If YES: classify into one of the 17 SoccerNet-V2 classes (Fig. 2) or label as “out-of-world” when relevant but off-taxonomy.
Intuition For many soccer matches, emotional moments such as goals and fouls are often accompanied by an excited tone in the commentary. This emotional context is a strong signal for the excitement judge, which can detect such moments without needing to rely on visual cues. The strategic context provided by the tactical judge is similarly informed by phrases that suggest changes in team strategy or formation.
Figure 2: A detailed look into the 17 action classes for Soccernet-v2.
Majority Voting System.
Once each judge makes its decision, we aggregate the results using a majority voting mechanism. If at least two of the three judges agree on the presence of a relevant action, the action is considered "spottable" and is classified as an event worthy of attention. If the judges disagree, the action is not classified as relevant. This ensures that only the most unanimously recognized actions are selected.
Out-of-World Action Classification.
In addition to the 17 predefined action classes, our system is designed to handle "out-of-world" actions—those that may be noteworthy but do not fall under any of the predefined classes. For instance, a player might execute a spectacular skill move or a controversial non-foul action, which can be exciting and relevant but doesn’t match the typical goal or penalty. In this case, the judges are given the opportunity to classify the action as "out of world," providing a broader view of game dynamics that goes beyond standard categories.
4 Evaluation and Results
We evaluate the performance of our proposed method using two metrics: average mean Average Precision (mAP) and tight mAP. The results on Soccernet-v2 Deliege et al. (2021) are summarized in Table 1, show that our LLM-based method achieves a significant improvement over existing baselines. Specifically, our method achieves an average mAP of 64.50% and a tight average mAP of 60.75%, outperforming the SoccerNet baseline Cioppa et al. (2020) by a margin of 14.80% in average mAP and 15.75% in tight mAP. It also surpasses RMS-Net Tomei et al. (2021) by 11.10% in average mAP and 11.25% in tight mAP.
Commentary-Anchored Random Baseline.
To establish a strict lower bound we create a commentary-anchored randomness baseline that predicts actions without reading the text. For each commentary sentence $s_{k}=(\tau_{k},\ell_{k},\text{text}_{k})$ in a half we sample a Bernoulli coin for every action class $c$ with probability equal to that class’s empirical commentary prior $p_{c}$ (estimated on the train split). If the coin succeeds we emit a pseudo detection $(\tau_{k},c,0.5)$; overlapping detections of the same class within $2\delta$ ($\delta{=}10$ s) are merged by keeping the earliest. This design respects the real timestamp distribution yet ignores all lexical information, yielding the hardest chance-level floor against which any text-aware model must improve.
Results.
When compared to the state-of-the-art method, ASTRA Xarles et al. (2023), our method achieves 96.51% of the average mAP and 97.23% of the tight average mAP, bringing us very close to the top-performing model. This demonstrates that our approach is highly competitive, offering strong performance with a focus on leveraging text-based information. Our method was implemented on an RTX 3060 GPU.
Method mAP (%) Tight
mAP (%)
CALF Cioppa et al. (2020) 49.7
RMS-Net Tomei et al. (2021) 53.4
ASTRA Xarles et al. (2023) 79.2 70.1
Random Text-Only 12.0 10.5
LLM-Based (Ours) 64.5 60.8
Table 1: Comparison of our method with existing approaches. Note that both metrics refer to average mAP as discussed in the beginning of Section 4. Our text-only method comes close to parameter-heavy highly trained action spotters.
Discussion.
Our current work raises the question if video frames, which are resource demanding in processing, needs to be included for action spotting. One can posit that this extends to a higher, more inclusive question: Do we need VLMs to spot any action where it is already being described in expert depth? Our current scope and focus in the paper lies in showing its capabilities for soccer match videos, which de-facto provide for expert-level realtime descriptions of events ongoing in the match. We notice compared to a blind lower bound baseline and the current SOTA VLM for action spotting, harnessing the abilities of few-shot LLMs is enough given the quality of commentary is rich. We plan to extend this to the aforementioned greater RQ and also provide a detailed study as to how quality of commentary affects any activity-based video action spotting based on our framework.
5 Limitations
While our framework shows promising results, there are several limitations to consider. First, the performance of our system is heavily dependent on the quality of the commentary and transcription. Inaccurate or incomplete commentary can hinder the ability of our judges to correctly identify action-worthy events, leading to lower accuracy in the action spotting task. Similarly, the quality of transcription performed by Whisper plays a critical role. Errors in the transcription process can result in incorrect words or misplaced timestamps, directly affecting the action spotting metrics, including mean Average Precision (mAP). These transcription errors could affect the reliability of the timestamped actions and ultimately influence the results of the semantic judging. Additionally, our framework assumes that the provided commentary is sufficiently detailed and relevant for the action spotting task. In cases where the commentary lacks context or important details, the system’s performance may degrade. We aim to address this in our future work.